if 2a+3b=9and 3a+2b=16 find the value of 3a-2b
Answer by mananth(16949)

2a+3b=9....1
3a+2b=16....2
multiply 1 by 3 and 2 by -2 and add the equations
6a+9b+-6a-4b=27-32
5b=-5
b=-1
plug the value of b in equation 1
2a-3=9
2a=6
a=3
3a-2b= 3*3 - 2*-1
=9+3=12
..